# Cluster-Norm for Unsupervised Probing of Knowledge

### v1

@@ Abstract
The deployment of language models brings challenges in generating reliable information, especially when these models are fine-tuned using human preferences. To extract encoded knowledge without (potentially) biased human labels, unsupervised probing techniques like Contrast-Consistent Search (CCS) have been developed (Burns et al., 2022). However, salient but unrelated features in a given dataset can mislead these probes (Farquhar et al., 2023). Addressing this, we propose a cluster normalization method to minimize the impact of such features by clustering and normalizing activations of contrast pairs before applying unsupervised probing techniques. While this approach does not address the issue of differentiating between knowledge in general and simulated knowledge—a major issue in the literature of latent knowledge elicitation (Christiano et al., 2021)—it significantly improves the ability of unsupervised probes to identify the intended knowledge amidst distractions.  [^1] The code for this work is available at https://github.com/Cadenza-Labs/cluster-normalization.
Machine Learning, ICML
@@ 1 Introduction
Figure 1: Under the standard CCS approach (left; Burns-Norm), a modified prompt including distracting random words causes all CCS probes to achieve random accuracy against ground truth labels (GT). When using our approach of cluster normalization (right; Cluster-Norm), the average accuracy of CCS probes increases significantly.
@@
Empirical evidence suggests that simulated internal beliefs or knowledge can be extracted from language model activations (Li et al., 2022; Gurnee & Tegmark, 2023; Azaria & Mitchell, 2023; Bubeck et al., 2023). Supervised probing methods can be employed to extract this knowledge (Alain & Bengio, 2016; Marks & Tegmark, 2023) but such methods require labels, which in some domains may not be provided correctly due to human biases or because humans simply do not know the correct label. It may even be critical to avoid the use of human labels to differentiate between a model’s true knowledge and its representation of human knowledge. Motivated by these ideas, unsupervised probing techniques like Contrast-Consistent Search (CCS) have been developed to extract the knowledge embedded in a language model without the need for ground truth labels (Zou et al., 2023; Burns et al., 2022).
Farquhar et al. (2023) outline current limitations of these approaches, demonstrating that these unsupervised probes tend to identify the most salient binary feature, which may not always correspond to the specific knowledge feature we seek. For example, in one experiment, one of a pair of distracting random words is added to each prompt in a text dataset. After training, unsupervised CCS probes function as classifiers for these random words, rather than the intended knowledge feature of the text. In practice, there may be numerous salient features of which we are unaware, which can divert an unsupervised probe from identifying the target feature $\mathcal{F}$, regardless of whether they are correlated or uncorrelated with $\mathcal{F}$.
To tackle this issue, we propose a cluster normalization method. Our method follows the usual initial approach of unsupervised probing of harvesting contrast pair activations, however we then cluster similar activations and normalize them separately, thereby eliminating the effect of distracting salient features. We can then apply any unsupervised probing method, such as CCS or CRC-TPC (Burns et al., 2022), to train a probe on these normalized activations. It is of course crucial to ensure this approach does not inadvertently eliminate the knowledge feature itself. To prevent this, we utilize contrast pairs, performing the clustering based on the average embedding of each pair. Further details on contrast pairs are provided in Section 2.1.
Probes trained with the original CCS approach achieve an average accuracy of approximately 0.5 on prompt datasets with distracting random word features (Farquhar et al., 2023). In contrast, our clustering method significantly improves this average accuracy to about 0.77, and to 0.81 for CRC-TPC when using Mistral-7b. Details of our method are described in Section 3.
@@ 2.1 Contrast-Consistent Search (CCS)
At first, a language model $\mathcal{M}$ processes a dataset of textual contrast pairs ${(x_{i}^{+},x_{i}^{-})}_{i=1}^{n}$, generating contextualized embeddings ${(\mathcal{M}(x_{i}^{+}),\mathcal{M}(x_{i}^{-}))}$. Following this, a linear probe (Alain & Bengio, 2016) is trained to calculate from these embeddings the probabilities ${p^{+}}$ and ${p^{-}}$, whether ${x_{i}^{+}}$ or ${x_{i}^{-}}$ is true, respectively. The objective function used to train this probe is given by a sum of two terms:
@@
Intuitively, there are at least two possible directions (features) satisfying this loss. The first is the knowledge direction we seek, and the second is the syntactical difference between positive and negative prompt templates. To remove this latter undesired feature, Burns et al. (2022) proceed as follows. Before training an unsupervised probe, contrast pair activations are first normalized: ~$\widetilde{\mathcal{M}}(x_{i}^{+})=\frac{\mathcal{M}(x_{i}^{+})-\mu^{+}}{\sigma^{+}}$, with $\mu^{+}$ and $\sigma^{+}$ the mean and standard deviation of the activations of all positive examples in each contrast pair; the same normalization procedure is followed for negative examples, and the unsupervised probe is trained on these normalized ~$\widetilde{\mathcal{M}}(x_{i}^{\pm})$. In this way, CCS removes the most salient feature of the contrast pair differences: the syntactical difference direction $\mathcal{F}_{\pm}$. However, as Farquhar et al. (2023) show, the second-most salient feature may not necessarily be the desired knowledge - an implicit assumption of the original CCS method. In this work, we take advantage of normalization to remove other undesired salient features by including a clustering step.
@@ 2.2 Contrastive Representation Clustering
As an alternative to CCS, the method of Contrastive Representation Clustering via Top Principal Component (CRC-TPC) (Burns et al., 2022) separates the normalized contrast pair differences {~$\{\widetilde{\mathcal{M}}(x_{i}^{+})-\widetilde{\mathcal{M}}(x_{i}^{-})\}$ based on projections onto their top principal component, i.e., the singular vector associated with the highest singular value, or the direction with the highest variance. This is again motivated by the intuition that the second-most salient contrastive feature after $\mathcal{F}_{\pm}$ removed by normalization - should be the knowledge feature $\mathcal{F}_{\top/\bot}$.
@@ 2.3 Theoretical Background
We shall first examine why there is a close link between the CCS loss described in Section 2.1 and the idea of saliency i.e., variance. Contrastive features will naturally achieve a low CCS loss. To see this, consider the variance of contrast pair differences projected along the feature direction of a given feature $\mathcal{F}$:
$X$ $:=\mathcal{F}^{T}\cdot\widetilde{\mathcal{M}}(x_{i}^{+}),~{}Y:=\mathcal{F}^{T}\cdot\widetilde{\mathcal{M}}(x_{i}^{-}).$
@@
In this expanded form, we see that the variance of contrast pair differences in the direction $\mathcal{F}$ captures,
  • confidence, with $E(X^{2})+E(Y^{2})$ higher if the magnitude projection along $\mathcal{F}$ in either element of a pair is high,
  • and consistency, with $-2\cdot E(X\cdot Y)>0$ if the projections of a contrast pair along $\mathcal{F}$ have opposing sign. This also increases with the magnitude of these projections.
@@
Due to this link between confidence, consistency, and saliency, a probe trained using the CCS loss will favor learning salient contrastive features. Otherwise, the projections of a contrast pair onto a feature will be small in difference or equal, failing to satisfy at least the consistency condition.
@@
  • $\mathcal{F}_{\pm}:=\mathcal{F}_{+}~{}-~{}\mathcal{F}_{-}$, the syntactical difference between contrast pairs due to the appending of positive and negative tokens, removed by normalization in the original CCS method,
  • $\mathcal{F}_{\top/\bot}:=\mathcal{F}_{\top}~{}-~{}\mathcal{F}_{\bot}$, the knowledge feature we seek.
@@
Let $f$ be some binary function, say the $XOR$ function on the presence of features, and $\mathcal{F}_{1}$, $\mathcal{F}_{2}$ be any two features. Suppose that a model represents the feature $f(\mathcal{F}_{1},\mathcal{F}_{2})$ as its own direction $F_{f(\mathcal{F}_{1},\mathcal{F}_{2})}$, orthogonal to $\mathcal{F}_{1,2}$ [^2] Evidence of such behavior has been observed with $f=XOR$ and any two features, as shown in (Marks, 2024).. Now, suppose both $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$ are distracting features and assume without loss of generality that exactly one of them appears in each pair with probability $\frac{1}{2}$.
@@
$\mathcal{M}(x_{i}^{+})-$ $\mathcal{M}(x_{i}^{-})=\underbrace{\mathcal{F}_{+}-\mathcal{F}_{-}}_{\mathcal{F}_{\pm}}+\underbrace{\mathcal{F}_{f(\mathcal{F}_{+},\mathcal{F}_{j})}-\mathcal{F}_{f(\mathcal{F}_{-},\mathcal{F}_{j})}}_{\Delta_{\mathcal{F}_{j}}^{\pm}}$
±(/)±$\pm(\underbrace{\mathcal{F}_{\top}-\mathcal{F}_{\bot}}_{\mathcal{F}_{\top/\bot}})\pm(\underbrace{\mathcal{F}_{f(\mathcal{F}_{\top},\mathcal{F}_{j})}-\mathcal{F}_{f(\mathcal{F}_{\bot},\mathcal{F}_{j})}}_{\Delta_{\mathcal{F}_{j}}^{\top}})$
@@
$E(\mathcal{M}(x_{i}^{+})-\mathcal{M}(x_{i}^{-}))=\mathcal{F}_{\pm}+\frac{1}{2}(\Delta_{\mathcal{F}_{1}}^{\pm}+\Delta_{\mathcal{F}_{2}}^{\pm})$ (1)
since $\mathcal{F}_{\pm}$ is constant, $\Delta_{\mathcal{F}_{j}}$ are both constant on half of the dataset and the two knowledge related terms have uniformly alternating sign. After centering, we have:
~$\widetilde{\mathcal{M}}(x_{i}^{+})-\widetilde{\mathcal{M}}(x_{i}^{-})=$/±$\pm\mathcal{F}_{\top/\bot}\pm\Delta_{\mathcal{F}_{j}}^{\top}$
$+\alpha\frac{1}{2}(\Delta_{\mathcal{F}_{1}}^{\pm}-\Delta_{\mathcal{F}_{2}}^{\pm})$
@@ 3 Method
Following the notation in Section 2.3, if $x_{i}$ belongs to cluster $c\in\{0,1\}$, a successful cluster normalization will leave:
~$\widetilde{\mathcal{M}}(x_{i}^{+})-\widetilde{\mathcal{M}}(x_{i}^{-})=\pm\mathcal{F}_{\top/\bot}\pm\Delta_{\mathcal{F}_{c}}^{\top}.$
@@
A key element to the effectiveness of our method is that our clustering approach does not erase the effect of the desired knowledge feature. This is achieved by clustering the averages of each contrast pair, $\mathcal{M}(x_{i})$. As a result, clustering only isolates salient non-contrastive features, and is effectively blind to $\mathcal{F}_{\pm}$ and $\mathcal{F}_{\top/\bot}$. Normalizing positive and negative samples separately per cluster aims to ensure that all contrastive features $\mathcal{F}^{\prime}$ related to $\mathcal{F}_{\pm}$ are properly normalized out - including the leaks from non-contrastive features $\mathcal{F}$ mixing with $\mathcal{F}_{\pm}$, as explained in Section 2.3. Note, we do not normalize out similar $\mathcal{F}^{\prime}$ resulting from the mixing of $\mathcal{F}$ with $\mathcal{F}_{\top/\bot}$. Eventually, only contrastive features related to knowledge are kept.
@@ 4 Experiments
In our experiments, we utilize Mistral-7B as our main language model, harvesting a ctivations (using the libraries from (Wolf et al., 2020) and (Nanda & Bloom, 2022)) at the 75th percentile layer (layer 24 for Mistral-7b) since, from preliminary experiments, we find probes achieve higher accuracies using the 50th to 90th percentile layers. We also report results using different language models (Phi-2 and 3, Gemma-7b, Llama-3-8B, Pythia-6.9b) and layers in the Appendices to verify the efficacy of our method. Our experiments follow the same general approach as those reported in Farquhar et al. (2023), as each of these original experiments set out to demonstrate the limitations of current unsupervised probing techniques.
@@
Activation clustering is performed using HDBScan, implemented in the scikit-learn library (Kramer & Kramer, 2016), setting a minimum number of elements in each cluster to 5 and using the Euclidean distance metric. One advantage of HDBScan over other clustering algorithms (e.g., k-means) is that the number of clusters does not need to be specified in advance. In order to examine the variance in probe performance, we report summary statistics of 50 probe fits in each experiment, and visualize the results from all.
The following experiments generally involve a comparison between an original prompt and a modified one, to attempt to induce a bias in an unsupervised probe. Hereon, we refer to these original prompts as unbiased and modified prompts as biased. We also refer to normalization over an entire dataset, as Burns normalization or Burns-Norm (See 2.3 for more details). We refer to our alternative approach through clustering as cluster normalization or Cluster-Norm. Unlike the approach in (Burns et al., 2022), where multiple prompt templates were used, the study in (Farquhar et al., 2023) utilized only one prompt template per dataset. Our method follows the prompt-template setup from (Farquhar et al., 2023).
Each experiment utilizes a train-test split of 70% for training and 30% for testing. Importantly, we evaluate our unsupervised probes on a test set where Burns-Norm is applied to the test set as it was done in (Burns et al., 2022), and not our cluster normalization. This is because we want probes to generalize, so if during evaluation they are fed with a contrast pair that belongs to an entirely different dataset, it is out of distribution for the clusters found during training. The probe should be a feature in the unaltered latent space. Although we do not use cluster norm on the test set for the aforementioned reason, we do use Burns-Norm for being able to compare our results with Burns et al. (2022) as this is what they do for the test set. Farquhar et al. (2023) likely follow a similar approach, as they mention utilizing normalization but do not provide any details regarding a train-test split.
For each experiment, we also report results using the CRC-TPC method as an alternative unsupervised probing technique to CCS. Finally, we report an upper-bound by using the results of the supervised method of logistic regression, similar as done by (Burns et al., 2022) and later also by (Farquhar et al., 2023).

### v2

@@ Abstract
The deployment of language models brings challenges in generating reliable information, especially when these models are fine-tuned using human preferences. To extract encoded knowledge without (potentially) biased human labels, unsupervised probing techniques like Contrast-Consistent Search (CCS) have been developed (Burns et al., 2022). However, salient but unrelated features in a given dataset can mislead these probes (Farquhar et al., 2023). Addressing this, we propose a cluster normalization method to minimize the impact of such features by clustering and normalizing activations of contrast pairs before applying unsupervised probing techniques. While this approach does not address the issue of differentiating between knowledge in general and simulated knowledge—a major issue in the literature of latent knowledge elicitation (Christiano et al., 2021)—it significantly improves the ability of unsupervised probes to identify the intended knowledge amidst distractions. [^1] The code for this work is available at https://github.com/Cadenza-Labs/cluster-normalization..
Cluster-Norm for Unsupervised Probing of Knowledge
Walter Laurito1,2, *, Sharan Maiya2,3, *, Grégoire Dhimoïla4,2, *, Ho Wan Yeung5, Kaarel Hänni6 1FZI, 2Cadenza Labs, 3University of Cambridge, 4ENS Paris-Saclay, 5UC Berkeley, 6Caltech Correspondence: laurito@fzi.de
**footnotetext: These authors contributed equally to this work.
@@ 1 Introduction
Figure 1: Under the standard CCS approach (left; Burns-Norm), a modified prompt including distracting random words (red) causes all CCS probes to achieve random accuracy against ground truth labels (GT), and high accuracy against these random word labels (random). When using our approach of cluster normalization (right; Cluster-Norm), the average accuracy of CCS probes for the desired feature increases significantly.
@@
Empirical evidence suggests that simulated internal beliefs or knowledge can be extracted from language model activations (Li et al., 2022; Gurnee and Tegmark, 2023; Azaria and Mitchell, 2023; Bubeck et al., 2023). Supervised probing methods can be employed to extract this knowledge (Alain and Bengio, 2016; Marks and Tegmark, 2023) but such methods require labels, which in some domains may not be readily provided due to human biases or because humans simply do not know the correct label. It may even be critical to avoid the use of human labels to differentiate between a model’s true knowledge and its representation of human knowledge. Motivated by these ideas, unsupervised probing techniques like Contrast-Consistent Search (CCS) have been developed to extract the knowledge embedded in a language model without the need for ground truth labels (Zou et al., 2023; Burns et al., 2022).
Farquhar et al. (2023) outline current limitations of these approaches, demonstrating that these unsupervised probes tend to identify the most salient binary feature, which may not always correspond to the specific knowledge feature we seek. For example, in one experiment, one of a pair of distracting random words is added to each prompt in a text dataset. After training, unsupervised CCS probes often function as classifiers for these random words, rather than the intended knowledge feature. In practice, there may be numerous salient features of which we are unaware, which can divert an unsupervised probe from identifying the target feature, regardless of whether they are correlated or uncorrelated with the target.
To tackle this issue, we propose a cluster normalization method. Our method starts by following the usual initial approach of unsupervised probing of harvesting contrast pair activations. We then cluster similar activations and normalize them separately, thereby eliminating the effect of distracting salient features. We can then apply any unsupervised probing method, such as CCS or CRC-TPC (Burns et al., 2022), to train a probe on these normalized activations. It is of course crucial to ensure this approach does not inadvertently eliminate the knowledge feature itself. To prevent this, we utilize contrast pairs, performing the clustering on the average embedding of each pair. Further details on contrast pairs are provided in Section 2.1.
Probes trained with the original CCS approach achieve an average accuracy of approximately 0.5 on prompt datasets with distracting random word features (Farquhar et al., 2023). In contrast, our clustering method significantly improves this average accuracy to about 0.77, and to 0.81 for CRC-TPC when tested with Mistral-7B. Details of our method are described in Section 3.
@@ 2.1 Contrast-Consistent Search (CCS)
At first, a language model $\mathcal{M}$ processes a dataset of textual contrast pairs ${(x_{i}^{+},x_{i}^{-})}_{i=1}^{n}$, generating contextualized embeddings ${(\mathcal{M}(x_{i}^{+}),\mathcal{M}(x_{i}^{-}))}$. Following this, a linear probe (Alain and Bengio, 2016) is trained to calculate from these embeddings the probabilities ${p^{+}}$ and ${p^{-}}$, whether ${x_{i}^{+}}$ or ${x_{i}^{-}}$ is true, respectively. The objective function used to train this probe is given by a sum of two terms:
@@
Intuitively, there are at least two possible directions (features) satisfying this loss. The first is the knowledge direction we seek, $\vec{F}_{\top/\bot}$, and the second is the syntactical difference between positive and negative prompt templates, $\vec{F}_{\pm}$. To remove this latter undesired feature, Burns et al. (2022) proceed as follows. Before training an unsupervised probe, contrast pair activations are first normalized: ~$\widetilde{\mathcal{M}}(x_{i}^{+})=\frac{\mathcal{M}(x_{i}^{+})-\mu^{+}}{\sigma^{+}}$, with $\mu^{+}$ and $\sigma^{+}$ the mean and standard deviation of the activations of all positive examples in each contrast pair; the same normalization procedure is followed for negative examples, and the unsupervised probe is trained on these normalized ~$\widetilde{\mathcal{M}}(x_{i}^{\pm})$. In this way, CCS removes the most salient feature of the contrast pair differences: the syntactical difference direction $\vec{F}_{\pm}$. However, as Farquhar et al. (2023) show, the second-most salient feature may not necessarily be the desired knowledge - an implicit assumption of the original CCS method. In this work, we take advantage of normalization to remove other undesired salient features by including a clustering step.
@@ 2.2 Contrastive Representation Clustering
As an alternative to CCS, the method of Contrastive Representation Clustering via Top Principal Component (CRC-TPC) (Burns et al., 2022) separates the normalized contrast pair differences {~$\{\widetilde{\mathcal{M}}(x_{i}^{+})-\widetilde{\mathcal{M}}(x_{i}^{-})\}$ based on projections onto their top principal component, i.e., the singular vector associated with the highest singular value, or the direction with the highest variance. This is again motivated by the intuition that the most salient contrastive feature after $\vec{F}_{\pm}$ removed by normalization - should be the knowledge feature $\vec{F}_{\top/\bot}$.
@@ 2.3 Theoretical Background
We shall first examine why there is a close link between the CCS loss described in Section 2.1 and the idea of saliency i.e., variance. Contrastive features will naturally achieve a low CCS loss. To see this, consider the variance of contrast pair differences projected along the feature direction of a given feature $\vec{F}$:
$X$ $:=\vec{F}^{T}\cdot\widetilde{\mathcal{M}}(x_{i}^{+}),~{}Y:=\vec{F}^{T}\cdot\widetilde{\mathcal{M}}(x_{i}^{-}).$
@@
In this expanded form, we see that the variance of contrast pair differences in the direction $\vec{F}$ captures,
  • confidence, with $E(X^{2})+E(Y^{2})$ higher if the magnitude of the projection along $\vec{F}$ in either element of a pair is high,
  • and consistency, with $-2\cdot E(X\cdot Y)>0$ if the projections of a contrast pair along $\vec{F}$ have opposing sign. This also increases with the magnitude of these projections.
@@
Due to this link between confidence, consistency, and contrastive saliency, a probe trained using the CCS loss will favor learning salient contrastive features. Otherwise, the projections of a contrast pair onto a feature will be small in difference or equal, failing to satisfy at least the consistency condition.
@@
  • $\vec{F}_{\pm}:=\vec{F}_{+}~{}-~{}\vec{F}_{-}$, the syntactical difference between contrast pairs due to the appending of positive and negative tokens, removed by normalization in the original CCS method,
  • $\vec{F}_{\top/\bot}:=\vec{F}_{\top}~{}-~{}\vec{F}_{\bot}$, the knowledge feature we seek.
@@
Let $f$ be some binary function, say the $XOR$ function on the presence of features, and $\vec{F}_{1}$, $\vec{F}_{2}$ be any two features. Suppose that a model represents the feature $f(\vec{F}_{1},\vec{F}_{2})$ as its own direction $\vec{F}_{f(\vec{F}_{1},\vec{F}_{2})}$, orthogonal to $\vec{F}_{1,2}$ [^2] Evidence of such behavior has been observed with $f=XOR$ and any two features, as shown in Marks (2024).. Now, fix two features $\vec{F}_{1}$ and $\vec{F}_{2}$ and assume without loss of generality that exactly one of them appears in each pair with probability $\frac{1}{2}$.
@@
$\mathcal{M}(x_{i}^{+})-$ $\mathcal{M}(x_{i}^{-})=\underbrace{\vec{F}_{+}-\vec{F}_{-}}_{\vec{F}_{\pm}}+\underbrace{\vec{F}_{f(\vec{F}_{+},\vec{F}_{j})}-\vec{F}_{f(\vec{F}_{-},\vec{F}_{j})}}_{\Delta_{\vec{F}_{j}}^{\pm}}$
±$\pm(\underbrace{\vec{F}_{\top}-\vec{F}_{\bot}}_{\vec{F}_{\top/\bot}})\pm(\underbrace{\vec{F}_{f(\vec{F}_{\top},\vec{F}_{j})}-\vec{F}_{f(\vec{F}_{\bot},\vec{F}_{j})}}_{\Delta_{\vec{F}_{j}}^{\top}})$
@@
$E(\mathcal{M}(x_{i}^{+})-\mathcal{M}(x_{i}^{-}))=\vec{F}_{\pm}+\frac{1}{2}(\Delta_{\vec{F}_{1}}^{\pm}+\Delta_{\vec{F}_{2}}^{\pm})$ (1)
since $\vec{F}_{\pm}$ is constant, $\Delta_{\vec{F}_{j}}$ are both constant on half of the dataset and the two knowledge related terms have uniformly alternating sign. After centering, we have:
~$\widetilde{\mathcal{M}}(x_{i}^{+})-\widetilde{\mathcal{M}}(x_{i}^{-})=$$\pm\vec{F}_{\top/\bot}\pm\Delta_{\vec{F}_{j}}^{\top}$
$+\alpha\frac{1}{2}(\Delta_{\vec{F}_{1}}^{\pm}-\Delta_{\vec{F}_{2}}^{\pm})$
@@ 3 Method
Figure 2: Mean accuracy of Logistic Regression, CRC-TPC, and CCS probes across six different models for (top) unmodified prompts and (bottom) modified prompts with distracting random words. Especially for the modified prompts, unsupervised methods using our Cluster-Normalization consistently outperform standard Burns-Normalization across the 25th, 50th, and 75th percentile layers and the final layer.
@@
Following the notation in Section 2.3, if $x_{i}$ belongs to cluster $c\in\{1,2\}$, a successful cluster normalization will leave:
~$\widetilde{\mathcal{M}}(x_{i}^{+})-\widetilde{\mathcal{M}}(x_{i}^{-})=\pm\vec{F}_{\top/\bot}\pm\Delta_{\vec{F}_{c}}^{\top}.$
@@
A key element to the effectiveness of our method is that our clustering approach does not erase the effect of the desired knowledge feature. This is achieved by clustering the averages of each contrast pair, $\mathcal{M}(x_{i})$. As a result, clustering only isolates salient non-contrastive features, and is effectively blind to $\vec{F}_{\pm}$ and $\vec{F}_{\top/\bot}$. Normalizing positive and negative samples separately per cluster aims to ensure that all contrastive features $\vec{F}^{\prime}$ related to $\vec{F}_{\pm}$ are properly normalized out - including the leaks from non-contrastive features $\vec{F}$ mixing with $\vec{F}_{\pm}$, as explained in Section 2.3. Note, we do not normalize out similar $\vec{F}^{\prime}$ resulting from the mixing of $\vec{F}$ with $\vec{F}_{\top/\bot}$. Eventually, only contrastive features related to knowledge are kept.
@@ 4 Experiments
In our experiments, we utilize Mistral-7B as our main language model, harvesting activations (using the libraries from (Wolf et al., 2020) and (Nanda and Bloom, 2022)) at the 75th percentile layer (layer 24 for Mistral-7B) since, from preliminary experiments, we find probes achieve higher accuracies using the 50th to 90th percentile layers. We also report results using different language models (Phi-2 and 3, Gemma-7B, Llama-3-8B, Pythia-6.9B) and layers in the following subsections and Appendices to verify the efficacy of our method. Our experiments follow the same general approach as those reported in Farquhar et al. (2023), as each of these original experiments set out to demonstrate the limitations of current unsupervised probing techniques. Individual results for each model can be found in Appendix B.1.
@@
Activation clustering is performed using HDBScan, implemented in the scikit-learn library (Kramer and Kramer, 2016), setting a minimum number of elements in each cluster to 5 and using the Euclidean distance metric. One advantage of HDBScan over other clustering algorithms (e.g., k-means) is that the number of clusters does not need to be specified in advance. In order to examine the variance in probe performance, we report summary statistics of 50 probe fits in each experiment, and visualize the results from all.
The following experiments generally involve a comparison between an original prompt and a modified one, to attempt to induce a bias in an unsupervised probe. Hereon, we refer to these original prompts as unbiased or non-biased and modified prompts as biased. We also refer to normalization over an entire dataset, as Burns normalization or Burns-Norm (See 2.3 for more details). We refer to our alternative approach through clustering as cluster normalization or Cluster-Norm. Unlike the approach in Burns et al. (2022), where multiple prompt templates were used, the study in Farquhar et al. (2023) utilized only one prompt template per dataset. Our method follows the prompt-template setup from Farquhar et al. (2023).
Each experiment utilizes a train-test split of 70% for training and 30% for testing. Importantly, we evaluate our unsupervised probes on a test set where Burns-Norm is applied to the test set as it was done in (Burns et al., 2022), and not our cluster normalization. This is because we want probes to generalize, so if during evaluation they are fed with a contrast pair that belongs to an entirely different dataset, it is out of distribution for the clusters found during training. The probe should be a feature in the unaltered latent space. Although we do not use cluster normalization on the test set for the aforementioned reason, we do use Burns-Norm for being able to compare our results with Burns et al. (2022) as this is what they do for the test set. Farquhar et al. (2023) likely follow a similar approach, as they mention utilizing normalization but do not provide any details regarding a train-test split.
For each experiment, we also report results using the CRC-TPC method as an alternative unsupervised probing technique to CCS. Finally, we report an upper-bound by using the results of the supervised method of logistic regression, similar as done by Burns et al. (2022) and later also by (Farquhar et al., 2023). For additional experiment details see also Appendix B.1.